OpenStudy (anonymous):

Calculus 3 Question: Write a parametric representation for the tangent line to the given curve at the given point in each of the following cases. (a) (cos 4t, sin 4t, t) at the point t = pi/8 (b) (t, 2t, t^2) at the point (1, 2, 1)

OpenStudy (anonymous):

anybody good with calculus 3 on openstudy?

OpenStudy (amistre64):

nobody that im aware of ...

OpenStudy (anonymous):

lol

OpenStudy (amistre64):

tangent to a curve is derivative of the curve function

OpenStudy (amistre64):

r = cos4t, sin4t, t r' = cos'4t, sin'4t, t'

OpenStudy (amistre64):

soo, x = cos4t + n (cos'4t) y = sin4t + n (sin'4t) z = t + n (t') all when t=pi/8

OpenStudy (amistre64):

and since thats in my own notative flair, you might have questions

OpenStudy (anonymous):

so first I get the derivative and then just plug in t?

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

and since having a scalar named "t" would have been confusing; i renamed it "n"

OpenStudy (anonymous):

so then I get r'(pi/8)=<4sin4(pi/8), -4cos4(pi/8), 1>

OpenStudy (amistre64):

cos' = -sin

OpenStudy (anonymous):

oh woops

OpenStudy (anonymous):

okay, the book says the answer is. <0, 1, pi/8>

OpenStudy (amistre64):

sin'=cos; so other than a few sign swaps; it looks fine

OpenStudy (amistre64):

for point or vector?

OpenStudy (amistre64):

your point ends in pi/8

OpenStudy (amistre64):

your vector ends in 1

OpenStudy (anonymous):

the "parametric representaion" is <0, 1, pi/8>

OpenStudy (anonymous):

^the answer

OpenStudy (amistre64):

r = cos4t, sin4t, t ; t=pi/8 r = 0,1, pi/8 is the point that we attach the tangent to

OpenStudy (amistre64):

the equation for the tangent line requires the tangent vector at that point to define it

OpenStudy (amistre64):

<0,1,pi/8> is not a parametrically defined tangent line; its simply a vector pointing to a point

OpenStudy (anonymous):

waiiit. I made a mistake, i copied the answer wrong

OpenStudy (amistre64):



OpenStudy (anonymous):

answer: (0, 1, pi/8) + t(-4, 0, 1)

OpenStudy (amistre64):

thats better

OpenStudy (amistre64):

but it makes no sense to say t = pi/8 and then put a nonrelated "t" into teh equation as a scalar

OpenStudy (amistre64):

AND ... that is not parametric form; that is vector form i believe

OpenStudy (anonymous):

okay. if you do r'(pi/8) you get (-4, 0, pi/8)

OpenStudy (amistre64):

t' = ?

OpenStudy (anonymous):

you don't need t' i don't think

OpenStudy (amistre64):

yes, you do, you need ALL the component derivatives

OpenStudy (anonymous):

t' would just be zero anyway

OpenStudy (amistre64):

t' = ?

OpenStudy (anonymous):

t = pi/8 t' = 0

OpenStudy (amistre64):

t = t t' = t' = 1

OpenStudy (amistre64):

after you get your derivative; then apply t=pi/8 to get a value

OpenStudy (amistre64):

pi/8' = 0 t' not= pi/8'

OpenStudy (anonymous):

okay I think I figured it out. first simply plug pi/8 into the line equation and you get r=<cos 4t, sin 4t, t> r(pi/8) = <0, 1, pi/8>

OpenStudy (anonymous):

next, plus pi/8 into the derivative of the line r'=<-4sin4t, 4cos4t, 1> r'(pi/8) = -4, 0, 1

OpenStudy (amistre64):

x = cos4t + n (cos'4t) y = sin4t + n (sin'4t) z = t + n (t') x = cos4t + n (-4sin4t) y = sin4t + n (cos4t) z = t + n (1) t = pi/8 ..................... x = 0 + n (-4) y = 1 + n (0) z = pi/8 + n (1)

OpenStudy (anonymous):

put it into vector form

OpenStudy (amistre64):

the question asks for parametric form; why would we need to put it into vector form?

OpenStudy (anonymous):

get the norm of r'(pi/8) which is <-4, 0, 1>, then use X + Nt = L (0, 1, pi/8) + t(-4, 0, 1)

OpenStudy (amistre64):

get the norm?

OpenStudy (amistre64):

the normal, if i read you right, is not the tangent ....

OpenStudy (anonymous):

sorry not the norm a parallel vector